The t value for a 98% confidence interval estimation with 24 degrees of freedom
(not the sample size n) is

1.317836

1.710882

2.063899

2.492159

2.796939
Answer #1

Solution :

Given that,

Degrees of freedom = df =24

At 98% confidence level the t is ,

= 1 - 98% = 1 - 0.98 = 0.02

/ 2 = 0.02 / 2 = 0.01

t /2,df = t0.01,24 = 2.492159

t value = 2.492159